OpenStudy (rockhauler98):

Write a rule for the sequence.(1 point) 3,-3,-9,-15,. . . (A)Start with 3 and add -6 repeatedly. (B)Start with -6 and add 3 repeatedly (C)Start with 3 and add 6 repeatedly (D)Start with 3 and subtract -6 repeatedly

rebeccaxhawaii (rebeccaxhawaii):

what is the first number for the sequence?

OpenStudy (codingking03):

A confederate flag? really? is that racist or what?

OpenStudy (rockhauler98):

no its not racist-_- and the answer to that is (C)

OpenStudy (rockhauler98):

am i correct

OpenStudy (amorfide):

you started with the number 3 your next term is -3 next term is -9 then next is -15 how do you get from 3 to -3 how do you get from -3 to -9

OpenStudy (rockhauler98):

idk im confused and i got the answer it was (C) and thxs i will medal

OpenStudy (amorfide):

you have 3 as your first number you want to get from 3 to -3 at this point it is either multiply by -1, or you add -6 so we look at the next term next term is -9 to get from -3 to -9 we either multiply by 3 or add -6 so the only thing in common is adding -6 therefore you start with 3 and add -6

OpenStudy (rockhauler98):

thank you

OpenStudy (amorfide):

the real way to do it

OpenStudy (amorfide):

arithmetic sequence is this a, a+d, a+2d, a+3d... so to work out if there is a common difference you would do 2nd term subtract the first term 3rd term subtract the second term 4th term subtract the third term which gives a+d - a= d a + 2d- a+d= d etc in this case -3 - 3=-6 -9--3=-6 -15--9=-6 this gives the difference of -6

OpenStudy (rockhauler98):

thxs

OpenStudy (amorfide):

for a geometric series you have \[a, ar, ar^{2}...ar^{n-1}\] and you would do 2nd term divided by first term 3rd term divided by second term 4th term divided by third term to work out the common ratio if they are the same answer it would be geometric

OpenStudy (rockhauler98):

okay okay i understand take a breath thank you;)